# A semi-supervised framework for diverse multiple hypothesis testing scenarios

### v2

@@ 7 Discussion
Each of the 10 datasets from Zhang et al. were downloaded from https://github.com/patrickrchao/AdaPTGMM_Experiments/tree/main/AdaFDR_experiments/data_files, containing the p-value and side information pairs, $(p,\mathbf{x})$. All pairs were then used to report a list of discoveries by each p-value based method with the following exceptions. In the RNA-seq datasets, Airway, Bottomly and Pasilla, there exists ‘spikes’ of high p-values. These ‘spikes’ concentrate in certain regions of the side information, i.e., when the log-normalized gene counts are low. Hence we removed those p-values with log-normalized gene counts that were less than zero. Figure S1 show the histograms of the p-values before and after this removal. The gene-drug response data was obtained from the adaptMT package.
@@
Figure S1: Histograms of p-values in the RNA-seq data before and after p-value removal. Top row: the original histograms of the p-values. Bottom row: the corresponding histograms after removal of the p-values associated with low counts.
S.17 Tables
Table S6: The number of discoveries and computations times for 13 of the PRIDE-20 dataset at the 1% FDR level. We calculated the number of discoveries at the 1% FDR using each method of Adaptive Knockoffs and TDC. We also computed computation times for Adaptive Knockoffs (in minutes). We used both the actual times and the estimated times according to Section S.13.
Project ID Method Estimated Time Actual Time Discoveries
PXD002470 AdaKO EM 0.58 1.17 300
AdaKO GAM 0.03 0.08 272
AdaKO LR 0.07 0.14 483
AdaKO RF 7.50 9.04 488
TDC - - 455
PXD006856 AdaKO EM 24.72 59.80 0
AdaKO GAM 0.80 2.17 379
AdaKO LR 3.67 4.63 428
AdaKO RF 427.01 486.70 421
TDC - - 425
PXD008920 AdaKO EM 33.81 48.63 8485
AdaKO GAM 0.95 2.39 8443
AdaKO LR 2.98 6.11 9143
AdaKO RF 250.66 270.67 9352
TDC - - 9086
PXD008996 AdaKO EM 119.73 188.54 7464
AdaKO GAM 3.39 7.55 7958
AdaKO LR 18.77 20.18 8481
AdaKO RF 1430.16 1626.96 8625
TDC - - 8413
PXD010504 AdaKO EM 134.07 185.59 459
AdaKO GAM 4.03 5.11 1358
AdaKO LR 9.25 11.37 1585
AdaKO RF 2134.38 2420.31 1626
TDC - - 1551
PXD012528 AdaKO EM 202.99 241.66 0
AdaKO GAM 6.37 8.34 239
AdaKO LR 13.57 16.34 314
AdaKO RF 3769.54 4143.21 405
TDC - - 342
PXD012611 AdaKO EM 4.79 7.59 2850
AdaKO GAM 0.15 0.26 3152
AdaKO LR 0.62 0.98 3250
AdaKO RF 43.53 47.01 3262
TDC - - 3260
PXD013274 AdaKO EM 34.65 49.52 14025
AdaKO GAM 0.81 1.70 14271
AdaKO LR 4.59 3.92 14791
AdaKO RF 205.10 248.82 14765
TDC - - 14769.0
PXD016724 AdaKO EM 499.32 598.31 3617
AdaKO GAM 13.76 24.01 3802
AdaKO LR 36.84 58.90 4536
AdaKO RF 8070.09 8882.65 4809
TDC - - 4745
PXD019186 AdaKO EM 247.15 389.51 0
AdaKO GAM 7.04 13.49 1020
AdaKO LR 34.77 30.35 942
AdaKO RF 4143.83 4427.20 1217
TDC - - 912
PXD022257 AdaKO EM 151.82 244.22 1619
AdaKO GAM 15.50 9.94 1918
AdaKO LR 12.34 23.81 2500
AdaKO RF 2475.32 2638.35 2501
TDC - - 2509
PXD023571 AdaKO EM 30.23 43.86 1030
AdaKO GAM 3.04 4.40 1743
AdaKO LR 2.92 10.34 0
AdaKO RF 437.75 474.17 2161
TDC - - 1505
PXD026895 AdaKO EM 62.89 85.62 3255
AdaKO GAM 8.94 4.94 3676
AdaKO LR 10.02 12.04 4044
AdaKO RF 847.23 835.02 4047
TDC - - 3980.0
S.18 Figures
Figure S2: Estimated FDR of each method in numerical simulations. Each panel plots the estimated FDR for each method at FDR thresholds ranging from 5% to 30%. The first row corresponds to Simulation 1 with three values of $k\in\{50,150,300\}$, the second row corresponds to Simulation 2 with three values of $k\in\{150,300,450\}$, and the last row corresponds to Simulation 3 and 4. For readability, the points are jittered in the horizontal direction.
Figure S3: Estimated power of each method in numerical simulations. Each panel plots the power for each method at FDR thresholds ranging from 5% to 30%. The first row corresponds to Simulation 1 with three values of $k\in\{50,150,300\}$, the second row corresponds to Simulation 2 with three values of $k\in\{150,300,450\}$, and the last row corresponds to Simulation 3 and 4. For readability, the points are jittered in the horizontal direction. A description of the AdaKO methods and KO can be found in Section 2.

### v3

@@ D.4 Preparation of data with p-values
All pairs of p-values and side information from the above datasets were used to report a list of discoveries using each p-value based method with the following exceptions. In the RNA-seq datasets, Airway, Bottomly and Pasilla, there exists ‘spikes’ of high p-values (Figure D1, top row). These spikes are associated with genes with low log-normalized read count. Hence, removing all genes with log-normalized read count less than zero — whose information content is questionable anyhow — also eliminated those undesirable spikes (Figure D1 bottom row). The number of genes removed is given in Table D5.
Dataset # of genes removed % of genes removed
Airway 10616 32%
Bottomly 2203 16%
Pasila 1575 13%
Table D5: The number and percentage of genes removed from the RNA-seq datasets.
@@
Figure D1: Histograms of p-values in the RNA-seq data before and after removal of genes with negative log-normalized read count. Top row: the original histograms of the p-values. Bottom row: the corresponding histograms after the removal of genes with low read counts.
Appendix E Supplementary results
E.1 Further comparisons between RESET and AdaKO in simulations
For each simulation and FDR threshold in Figure 1 of the main text, we determined the maximum power across all RESET and AdaKO methods and computed the relative power of each method to this maximum. The resulting relative powers are reported as boxplots in Figure E1A, giving an indication of the performance of each method across the simulations as a whole. Additional plots showing the performance of each method in each separate simulation are given in Figures E2 (power) and E3 (estimated FDR).
While in Figure E2 we often find a version of AdaKO that is more powerful, or at least comparable to RESET Ensemble, this varies between the versions. On the other hand, RESET Ensemble’s performance is consistent across these simulations, and overall typically achieves the highest relative power when the results are aggregated. Interestingly, RESET NN seems to be marginally more powerful than RESET Ensemble. While this may be the case, we still recommend RESET Ensemble as the default approach to alleviate the burden of selecting any single classification algorithm, as for example, RESET Ensemble appears to perform marginally better in the p-value setting (see Figure 6 of the main text).
Figure E1: Relative power and runtimes for each method. (A) Each boxplot describes the relative power for each method, computed across the combination of simulations and FDR thresholds in Figure 1. (B) The log10 runtimes of each method applied at the 5% FDR level for 5 runs of Simulation 4. For readability, the points are jittered in the horizontal direction.
One advantage of RESET is its computational speed. Indeed, our implementation of RESET, as described in Section 4.4 of the main text, has many steps that can be parallelized: the $r$ evaluations of each classification algorithm on the $K$ folds can all be done in parallel using multiple cores. In contrast, Adaptive Knockoffs must iteratively apply a classification algorithm for every nonzero scoring hypothesis in the candidate set $\mathcal{\hat{S}}_{t}$ until the estimated FDR is $\leq\alpha$, or all hypotheses are revealed. Figure E1B demonstrates this advantage showing the (log10) runtimes of RESET (using 20 CPUs) and of Adaptive Knockoffs at the 5% FDR level across 5 runs of Simulation 4. We find that each version of RESET is significantly faster than each version of Adaptive Knockoffs. RESET Ensemble is around $10^{3}$ to $10^{4}$ times faster than AdaKO EM, $10^{2}$ to $10^{3}$ times faster than AdaKO RF, and about 3-5 times faster than AdaKO LR and AdaKO GAM.
Figure E2: Estimated power of each method in numerical simulations. Each panel shows the power for each method at FDR thresholds ranging from 5% to 30%. The first row corresponds to Simulation 1 with three values of $k\in\{50,150,300\}$, the second row corresponds to Simulation 2 with three values of $k\in\{150,300,450\}$, and the last row corresponds to Simulations 3 and 4. For readability, the points are jittered in the horizontal direction. A description of the AdaKO methods and KO can be found in Section 2.
Figure E3: Estimated FDR of each method in numerical simulations. Each panel shows the estimated FDR for each method at FDR thresholds ranging from 5% to 30%. The estimated FDR is calculated as the proportion of irrelevant variables in the discovery list, averaged over the runs. The first row corresponds to Simulation 1 with three values of $k\in\{50,150,300\}$, the second row corresponds to Simulation 2 with three values of $k\in\{150,300,450\}$, and the last row corresponds to Simulations 3 and 4. For readability, the points are jittered in the horizontal direction.
E.2 Comparison of AdaKO RF and RESET Ensemble using PRIDE data
We looked at the performance of RESET Ensemble versus AdaKO RF because in this setup it was the most powerful implementation of Adaptive Knockoffs. Due to the computational constraints of AdaKO RF, we focused on the results using a narrow search. In Figure E4A, the left boxplot displays the average number of discoveries by RESET Ensemble, divided by the number of discoveries by AdaKO RF using thirteen of the PRIDE-20 spectrum files. In six of the thirteen PRIDE spectrum files, we found RESET Ensemble to be the superior method, while in the remaining seven spectrum files, AdaKO RF produced the same or more discoveries. In one dataset, AdaKO RF discovered 1,217 peptides — approximately an 11% increase over the average of 1,084 peptides discovered by RESET Ensemble.
Figure E4: Comparison of RESET Ensemble and AdaKO RF. For each of the 13 PRIDE datasets, we computed the average number of discoveries using 10 applications of RESET Ensemble, varying RESET’s internal seed. The left panel shows a boxplot of those RESET Ensemble averages divided by the number of discoveries using AdaKO RF (left boxplot) and TDC (right boxplot) where each circle corresponds to one of the 13 smaller PRIDE-20 datasets. In the centre panel, we compare the estimated and actual runtimes for AdaKO RF. Lastly, in the right panel, we compare the actual runtimes of RESET Ensemble and AdaKO RF (in log10 scale) on the same 13 PRIDE-20 datasets.
We also assessed the computational times of AdaKO RF in Figure E4B, using both the actual and estimated runtimes as outlined in Section C.14. Reassuringly, our estimated times were less than the actual times taken by AdaKO RF to complete. Although we find that, when analysing these smaller spectrum sets, AdaKO RF is the most powerful method of Adaptive Knockoffs, and it is competitive to RESET Ensemble power-wise, it is again clear that AdaKO RF is generally too slow with some runtimes exceeding multiple days even for these small datasets. In fact, on the dataset where AdaKO RF outperforms RESET Ensemble by about 11%, it takes 3.07 days for AdaKO RF to discover its peptides while RESET Ensemble takes about half a minute to complete. Figure E4C shows that in the worst cases, RESET takes roughly 1 minute while AdaKO RF takes over 6 days. The issue of runtimes is especially problematic considering the fact that multiple spectrum files are often analysed jointly, as in our HEK293 analysis in Section 5.2.1.
E.3 FDX control in mass spectrometry data
We evaluated RESET Ensemble’s FDX control using the HEK293 and all twenty PRIDE-20 spectrum files (see Sections D.1 and D.2 for data preparation and search settings). As in Section 5.2.1, the joint HEK293 spectrum files were searched against 5 distinct target-decoy databases, where each target-decoy database was generated by concatenating the target database to a decoy database consisting of shuffled target peptides. Similarly, each PRIDE-20 spectrum file was searched against 10 distinct target-decoy databases. We used the coinflip version of FDP-SD (Luo et al. (2023)), which provides a uniform increase in the number of discoveries (at the cost of some randomness in the discovery list, see Supplementary Algorithm A2). Our results using RESET Ensemble and FDP-SD are given in Table E1 and Figure E5.
Table E1: The average number of (HEK293) discoveries while controlling the FDX. The average number of discoveries reported for RESET Ensemble and FDP-SD on the combined HEK293 spectrum files at different confidence values $1-\gamma=0.5,0.8,0.9$ and an FDP threshold of 1%. The averages are taken over 5 applications of each method, once for each of the 5 combined target-decoy databases.
Confidence $1-\gamma$
Method 0.5 0.8 0.9
RESET Ensemble 82930 82677 82508
FDP-SD 75977 75450 75255
RESET Ensemble delivers on average approximately 9-10% more HEK293 discoveries than FDP-SD across all confidence levels (Table E1). The picture is somewhat murkier when considering the PRIDE-20 datasets: while RESET Ensemble appears to be generally more powerful (Figure E5, zoomed-in right panel), it fails to make any discoveries at high confidence levels in a few datasets (Figure E5, left panel). Upon a closer look, the numbers of discoveries made by FDP-SD in these datasets are relatively small, with the largest average number of discoveries among these datasets being 417. When the number of discoveries is expected to be small and the confidence parameter is high, RESET may yield less power, even if the side information is reasonably informative — a phenomenon we dwell on in Section F.4.
RESET Ensemble advantage over FDP-SD is clearer when applying the two procedures to open as opposed to narrow searches. As discussed in Section D.2, an open search allows for the detection of peptides whose masses have been altered by post-translational modifications (PTMs). In this setting, side information variables such as dm, which records the mass difference between the peptide and the spectrum, can be highly informative given that some differences will correspond to masses of common PTMs. Thus, it is not surprising that RESET, which can take advantage of such side information, improves on FDP-SD even further across all confidence levels in this context.
Figure E5: Comparison of RESET Ensemble and FDP-SD. The left panel shows boxplots of the average number of discoveries using RESET divided by the average number of discoveries using FDP-SD (the right panel is just a ‘zoomed-in’ version). We used all twenty PRIDE-20 datasets at an $\alpha=1\%$ FDP threshold and varying confidence parameters with both narrow and open search modes. The averages are taken over 10 applications, one for each of the 10 target-decoy database pairs we generated for each of the PRIDE-20 dataset.
E.4 Result-related figures
Figure E6: Comparison of RESET and other methods using a collection of publicly available datasets. Each method was evaluated at the FDR thresholds of 1%, 5%, 10%, and 20%. The number of discoveries reported by each RESET method is averaged over 10 applications at FDR threshold and dataset. Each colour and shape combination corresponds to a unique method. For readability, each point is jittered in the horizontal direction.
Appendix F Supplementary analyses
F.1 Analysis of AdaDetect and empirical p-values
The main assumption of AdaDetect’s FDR control in the novelty detection problem, is that the null observations $\left(Y_{1},\dots,Y_{n}\right)$ and the null test set $\left(X_{i}:X_{i}\sim P_{0}\right)$ are exchangeable conditioned on the novelties $\left(X_{i}:X_{i}\not\sim P_{0}\right)$ (Marandon et al. (2024)). In the multiple testing setup, the closest analogue to the null observations is the collection of decoy information $\{(\mathbf{x}_{i},W_{i}):L_{i}=-1\}$, and the analogue to the null test set is the true null target information $\{(\mathbf{x}_{i},W_{i}):L_{i}=1,i\in N\}$. However the exchangeability assumption of AdaDetect in this setting fails for two reasons: (1) the collection of decoy information $\{(W_{i},\mathbf{x}_{i}):L_{i}=-1\}$ contains a mixture of true and false null hypotheses which cannot be discerned and (2) the true null hypotheses can be heterogeneous.
Despite failing to meet the assumptions for guaranteed FDR control, we nonetheless verified this in practice through the following simulation inspired by Zhang et al. (2019). We generated $m=20K$ hypotheses with side information $\mathbf{x}_{i}\sim U[0,1]$, each drawn independently. Each hypothesis is declared a false null with probability $\pi_{i}$ and a true null otherwise, where
$\pi_{i}=\pi(\mathbf{x}_{i}):=0.1\cdot\varphi^{c}_{[0.75,0.05]}\left(\mathbf{x}_{i}\right),$
and $\varphi^{c}_{[\mu,\sigma]}$ denotes the 0-1 clamping (truncating between 0 and 1) of the normal density with mean $\mu$ and standard deviation $\sigma$. For each hypothesis we drew $z_{i}\sim\mathcal{N}(\mu_{i},1)$ where
[…]
where $B_{i}\in\{0,-0.5\}$ with equal probability and independent of everything else. One sided p-values are then constructed as $p_{i}=1-\Phi(z_{i})$, where $\Phi$ is the standard normal CDF. This construction allows us to create a heterogeneous mixture of true nulls that violate the exchangeability assumptions.
We evaluated two versions of RESET on the simulated data, the first is the standard RESET Ensemble as described in Section 4.4 in the main text, and the second adopts AdaDetect’s empirical p-value approach, which we now describe. For each target ($L_{j}=1$), we defined the empirical p-value by comparing the target’s score with the scores of the estimating decoys:
$p_{j}:=\frac{1+\#\{k:\widetilde{W}_{k}>\widetilde{W}_{j},\tilde{L}_{k}=+1,L_{k}=-1\}}{1+\ell},$ (8)
where $\ell:=\#\{k:\tilde{L}_{k}=+1,L_{k}=-1\}$ is the number of estimating decoys. Next, we used Storey’s procedure which estimates the fraction of true null hypotheses, $\pi_{0}$, and applies the BH procedure at the adjusted level of $\alpha/\hat{\pi}_{0}^{Storey}$. Storey’s estimate is given by
$\hat{\pi}_{0}^{Storey}:=\frac{1+H(\lambda)}{(1-\lambda)\cdot\#\{L_{k}=+1\}},$
where $H(\lambda)$ denotes the number of empirical p-values $\geq\lambda$. We used $\lambda=\frac{\lfloor\ell/2\rfloor}{\ell+1}$. This empirical p-value version of RESET is essentially applying the AdaDetect methodology. The only difference is that we can conduct a fair comparison in which both versions of RESET offer the same semi-supervised learning approach.
Figure F1 shows the empirical FDRs using the empirical p-value version of RESET Ensemble (denoted as RESET Ensemble-p) versus the usual RESET Ensemble, averaged over 100 realizations of the data. At 10% and 20% FDR thresholds, we find that empirical FDRs calculated for RESET Ensemble-p are at least 5 standard errors above their respective thresholds, empirically demonstrating failure to control the FDR. Emery et al. (2020) show a similar failure to control the FDR using BH or Storey’s procedure using pooled empirical p-values when the true null hypotheses are heterogenous (which they call “noncalibrated”).
Figure F1: Estimated FDRs with and without empirical p-values. The confidence intervals correspond to $\pm 5$ times the estimated standard error. For readability, points are jittered in the horizontal section. The horizontal segments mark the FDR thresholds on the y-axis.
F.2 Brief justification of the assumptions used with HEK293 and PRIDE-20 datasets
TDC (equivalently SSS+) has become the de facto method of FDR analysis in the proteomics literature since its introduction by Elias and Gygi (2007). An assumption of TDC for valid FDR control is that the label $L_{i}$, indicating whether an incorrect target or its decoy scored higher, is uniformly $\pm 1$ independently of all the scores as well as of the labels corresponding to the correct targets (Assumption 1 of the main text). This assumption relies heavily on the decoy construction which we now outline.
There are two main approaches to decoy construction: randomly shuffling or reversing each target peptide sequence. This random shuffling or reversing is typically applied to the non-terminal amino acids, i.e., the amino acids between the first and last. The rationale is that the terminal amino acids of target peptides are typically of a certain type, e.g., lysine or arginine, and therefore they can be used to distinguish target and decoy peptides if they are not preserved. By either shuffling or reversing in this way, the result is a set of decoy peptides that have the exact same length, amino acid profile, and terminal amino acids as their corresponding target peptides, thus creating realistic competition for incorrect target discoveries in an attempt to satisfy the above assumption. While both approaches are commonly used, in this paper we use randomly shuffled decoys so that our results can be averaged across different decoy generations.
Importantly, the above assumption has been corroborated empirically through entrapment experiments (Granholm et al. (2011)). In these experiments, a sample of known peptides is used so that the actual FDP can be gauged. This strategy has allowed others (He et al. (2015); Lin et al. (2022)) to show that the FDR is controlled using TDC, suggesting that Assumption 1 are reasonable in practice.
RESET additionally requires this statement to be true if we further condition on the side information $\mathbf{x}$. This is obviously true for side information variables that are constant between the target and corresponding decoys — e.g., PepLen, Charge, lnNumSP, dm, absdM. The other side information variables are score-related — e.g., Xcorr, deltCn and deltLCn, and so conditioning on them assumes no more than the ‘standard assumption’ that already conditions on such a score in the first place. Hence, we believe Assumption 3 with $P(L_{i}=1)=c_{0}=1/2$ to be reasonably satisfied. Moreover, in our recent publication, we empirically demonstrated through entrapment experiments that RESET (combined with Percolator, a semi-supervised algorithm designed for peptide detection) controls the FDR (Freestone et al. (2025)).
F.3 Analysis of Adipose datasets
Figure F2: Comparison of power in Adipose data. In (A) and (B), we compare the number of discoveries using the two eQTL datasets. The number of discoveries reported by each RESET method is averaged over 10 applications at a given FDR threshold and dataset. In both panels, we evaluated each method at FDR thresholds of: 1%, 5%, 10%, 20%.
Notably, we have excluded from Figure 6 the analysis of the two eQTL studies for the following reason. Due to linkage disequilibrium these datasets exhibit a high degree of p-value as well as side-information dependency, even among the true null hypotheses. Hence these datasets violate Assumption 3 which guarantees our — as well as the other considered methods’ — type-1 error control. It is thus unclear whether our comparisons would remain fair, particularly if one method violates FDR control more than the others. Notably, RESET Ensemble flagged this data as exhibiting suspicious dependence using the feature outlined in Section C.2.4.
Figure F2(A-B) shows that overall RESET Ensemble is able to report a similar number of discoveries to other methods, with the exception of $\alpha=20\%$ in the Adipose Subcutaneous dataset, where RESET Ensemble reports substantially less discoveries. Even if we overlook the dependency in the data, and include the two eQTL studies in our power analysis in Figure 6, RESET Ensemble still ranks the highest. Indeed, Figure F3 shows the analogous results with the eQTL studies included. RESET Ensemble has the highest median at 93.6%, followed by AdaFDR at 93.2% and RESET NN at 91.8%.
Figure F3: Relative power of each method. The boxplots show the relative power of each method, calculated for each dataset described in Section D.3 (including the two Adipose datasets) and across FDR thresholds of 1%, 5%, 10% and 20%.
F.4 Further analysis of RESET with FDX control
We point out the following limitation on the topic of RESET’s FDX control. When the number of discoveries is small, e.g., when the confidence is too high, RESET may yield less power, even if the side information is reasonably informative. To illustrate this phenomenon, we computed the bounds, $\delta_{i}$ on the number of decoy wins in the top $i$ scores from Algorithm A2, which are used to determine the reported list of discoveries in FDP-SD (with $c=1/2$) and RESET (using the default $c=2/3$). We considered a confidence level of $1-\gamma=90\%$ and an FDP threshold of $\alpha=1\%$. Since we removed approximately half the decoys for training purposes in RESET, the bounds need to be doubled to fairly compare between the two approaches.
Refer to caption Refer to caption
Figure F4: RESET’s and FDP-SD’s bounds on the number of decoy wins. We recorded, on the left y-axis, the bounds used to determine the list of FDP-SD discoveries and twice the bounds used to determine RESET’s discoveries at a range of indices at $\alpha=1\%$ and $1-\gamma=90\%$. On the right y-axis, we plot the ratio of these bounds (in black). The left panel looks at indices between 501 to 2K, while the right panel looks at indices from 10K to 20K. An index corresponds to the number of top scoring hypotheses (regardless of their labels).
Figure F4 shows the bounds for FDP-SD, and double the bounds for RESET, along with the ratio of the two at a range of different indices. We can see that for smaller indices, this ratio is much smaller, initially at 0% at an index of 501 and 50% at an index of 1K, while at 10K this ratio is much higher at 94%. If the order of the hypotheses is the same, then RESET’s smaller bounds imply it will report fewer discoveries because RESET employs FDP-SD to control the FDX by searching for the largest index $i$ s.t. $A_{i_{0}}\leq\delta_{i_{0}},\dots,A_{i}\leq\delta_{i}$, and reports the top $i$ scoring positively labeled hypotheses. In other words, if RESET is to report roughly the same number of discoveries as FDP-SD, then RESET needs to rearrange the targets and estimating decoys successfully enough so as to make up for a 50% difference in the bounds when considering the top 1K scores, while only 6% when considering the top 10K. If the side information is not rich enough, then it is possible RESET will report fewer discoveries than FDP-SD in these cases. We intend to investigate this as part of future work.
F.5 Variability of RESET
RESET’s discovery list is variable due to the random split of the decoys. This variability is demonstrated in the histogram of the number of discoveries in Figure F5A using 100 applications of RESET Ensemble to the Airway dataset at an FDR threshold of 10%. The resulting number of discoveries has a mean of 6018 with a standard deviation of about 75.
While this is undesirable, this analysis does not take into account the randomness of the data, and therefore that the added variance from RESET may be marginal. Indeed, in Figure F5B, we find that in the case of Simulation 5 from Section 6.2.1, RESET is overall arguably less variable than AdaPT, even though AdaPT does not employ any internal randomization. In practice, users may get around RESET’s randomization by setting the random number generator to be a function of the input data and parameters. We plan on adding this feature to a future update.
Figure F5: Variability in the number of discoveries. (A) Histogram of the number of RESET Ensemble discoveries applied 100 times to the Airway data with a 10% FDR threshold and varying its internal seed. (B) Boxplots of RESET’s and AdaPT’s power in 100 applications of the ‘circle in the middle’ setup of Simulation 5 using a 10% FDR threshold. There is one major outlier in RESET’s case (which occurred when the signal was considerably low).